OpenStudy (anonymous):

Anne Katz, owner of Katz Sport Shop, loans $8,000 to Shelley Slater to help her open an art shop. Shelley plans to repay Anne at the end of 8 years with interest compounded semiannually at 8 percent. At the end of 8 years, Anne will receive

OpenStudy (gw2011):

Use the following formula: A=P(1+i)^n A=The amount to be received P=Principal i=interest--in this case it will be0.08/2=0.04=4% because 8% is compounded semiannually n=Number of periods the interest applies which = 16 because 0.04 semiannually is 2 times per year and for 8 years this equals 16 periods A=(8000)(1.04)^16 A=(8000)(1.873) A=$14,984